OpenStudy (anonymous):

can someone explain me how to solve this problem 21 + 3x = 42

OpenStudy (anonymous):

subtract 21 on each side, and then divide by 3:)

OpenStudy (anonymous):

your answer would be x = 7

OpenStudy (pinksapphire):

21+3x=42 -21 -21 3x=42 3x/3=21/3 x=7

OpenStudy (anonymous):

so 21 - 21 + 3x = 42 - 21 and then you would get 0 + 3x divided by 3 = 21 divided by 3 right

OpenStudy (pinksapphire):

Yes :)

OpenStudy (anonymous):

no, you would get 3x = 21, and then divide each side by 3 to isolate the x and get x = 7 :)

OpenStudy (anonymous):

ok that's kelly234

OpenStudy (anonymous):

move 21 over, then its 21 since when you move you have to switch the sign. now you have 3x = 21. divide each side by 3 and then bam you get x = 7

OpenStudy (anonymous):

do u get what i did @AustinDaMan ?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

yw:) you should click on the "best response" by my name:)